OpenStudy (anonymous):

Find all pairs ( a, b ) such that 2a + b = 12 and ab = 3.

OpenStudy (zzr0ck3r):

a=3/b (b is not 0) so 2a+b=12 is the same as 2(3/b)+b = 12 6/b+b=12 6+b^2=12 b^2=6 b = sqrt(6) so b = 3/sqrt(6)

OpenStudy (zzr0ck3r):

+- sqrt(6)